OpenStudy (anonymous):

how do I solve -10(x+4)+48=-16-16

OpenStudy (nory):

We want to get the x on one side alone. First, the right side can be simplified. -16-16 = -32. Now we have -10(x+4) + 48 = -32. Subtract 48 from both sides to get -10(x+4) = -80. Now we can do a little trick: divide both sides by -10. x+4 = 8 x = 4 Does that make sense?

OpenStudy (texaschic101):

-10(x + 4) + 48 = -16-16 (distribute the -10 through the parenthesis) -10x - 40 + 48 = -16-16 (combine like terms) -10x + 8 = -32 (subtract 8 from both sides) -10x = -32 - 8 -10x = -40 (divide by -10 on both sides) x = -40/-10 (negative divided by a negative equals a positive) x = 4